OpenStudy (anonymous):

This is weird. Why is it that every time I square -1, I get -1 but when I do -1*-1 on my calculator it comes out as 1. Is there something wrong with my calculator.

OpenStudy (zzr0ck3r):

\[-1^2 = -(1)^2 = -(1) = -1\\\text{and}\\(-1)^2=(-1)(-1) = 1\]put this into your calculator \[(-1)^2\]

OpenStudy (zzr0ck3r):

in general \[-a^b \ne (-a)^b\]unless b is odd

OpenStudy (anonymous):

Oh I get it nm. Omg ur awesome!

OpenStudy (zzr0ck3r):

-a=-1*a

OpenStudy (zzr0ck3r):

np